# Misaligning Reasoning with Answers - A Framework for Assessing LLM CoT Robustness

### v1

@@ Abstract
LLMs’ decision-making process is opaque, prompting the need for explanation techniques like Chain-of-Thought. To investigate the relationship between answer and reasoning, we design a novel evaluation framework, MATCHA. In domains like education and healthcare, reasoning is key for model trustworthiness. MATCHA reveals that LLMs under input perturbations can give inconsistent/nonsensical reasoning. Additionally, we use LLM judges to assess reasoning robustness across models. Our results show that LLMs exhibit greater vulnerability to input perturbations for multi-step and commonsense tasks than compared to logical tasks. Also, we show non-trivial transfer rates of our successful examples to black-box models. Our evaluation framework helps to better understand LLM reasoning mechanisms and guides future models toward more robust and reasoning-driven architectures, enforcing answer-reasoning consistency.
@@ 1 Introduction
Large Language Models (LLMs) like GPT (gpt4, ), Llama (llama3, ), DeepSeek (deepseekr1, ) have demonstrated remarkable capabilities in complex reasoning tasks through techniques such as chain-of-thought (CoT) prompting (cot, ; wei2022chain, ; xu2024preemptive, ). By decomposing problems into intermediate reasoning steps, CoT enables LLMs to tackle mathematical proofs wang2025malotmultiagentleanbasedlong , symbolic manipulations xu-etal-2024-faithful , and multi-step inference wang-etal-2023-towards problems that standard prompting often fails to solve.
Despite the widespread adoption of Chain-of-Thought (CoT) prompting and the rapid integration of large language models (LLMs) across a variety of domains (hadi2023survey, ; filippo2024future, ), the mechanisms underlying CoT reasoning remain poorly understood (wang2022towards, ). Despite this gap in our understanding, LLM CoT is already being used in practical scenarios to provide explanations. For example, student’s often receive an answer key but can rely on LLMs to explain how they should arrive to these answers grassucci2025beyond . Additionally, medical patients can rely on LLM explanations to understand their diagnoses ali2024improving . Giving wrong explanations to students can further their misunderstandings and receiving wrong information on a diagnosis degrades trust in medicine medicaltrust . This raises a key question: can we trust CoT reasoning to be robust? As shown in the examples above, when reasoning is fragile it can contradict the answer, obscure the method, or even mislead entirely. Small changes to the input which have a small effect on the question should have little to no impact on the reasoning; however, as we will show in this paper, this is often not the case. In high-stakes settings, the right answer alone is not enough, stable and accurate reasoning is also essential for building usable real-world systems.
@@
In this paper, we show the fragility of LLM CoT, as exemplified in Figure 1, where a perturbed question leads to the wrong reasoning yet correct answer (DeepSeek-R1-7 on GSM8K). We adopt an answer-first, then reasoning framework to isolate failures in reasoning while keeping the answer fixed. This setup not only exposes reasoning-specific vulnerabilities but also reflects real-world scenarios where LLM reasoning is often generated post hoc to justify already-decided outputs jeyasothy2023general ; xu2024preemptive . To this end, we introduce MATCHA, a novel algorithm for finding small input changes that cause inconsistent behavior in CoT reasoning. We use MATCHA to create metrics for evaluating the robustness of an LLM’s CoT for different models. MATCHA can be instantiated as either a token-level or embedding-level perturbation. Our primary focus is, token-level perturbations as they are more realistic, however, they are challenging to create due to the need to introduce minimal, semantically plausible edits that selectively disrupt reasoning without altering answer correctness. To achieve this, we first randomly insert tokens into the original question. Then, we identify the inserted tokens that have the largest influence on the reasoning process while ensuring that answer-related tokens remain unchanged. By leveraging gradient-based importance ranking, we selectively replace tokens that maximize reasoning perturbation while minimizing changes to the original question. MATCHA as an embedding-level perturbation perturbs the embedding space of input questions imperceptibly, ensuring that the token sequence remains unchanged while disrupting the reasoning process. By optimizing a loss that maximizes reasoning divergence while preserving the correct answer, the perturbation subtly shifts the internal representations, leading to incorrect reasoning without altering the final answer. Our code is available at https://github.com/uiuc-focal-lab/MATCHA.
@@
  • We are the first to study the misalignment between answer and reasoning systematically. We introduce a novel evaluation framework MATCHA (Misaligned Answer and Thought CHAin) which works at both the token- and embedding-level.
@@ 2 Related Work
Prompt-based Reasoning. Chain-of-Thought was first introduced by cot , showing that inducing intermediate reasoning steps significantly improves the reasoning ability of LLMs, which is widely regarded as an important form of NLEs. Many works have further expanded on Chain-of-Thought (CoT) using self-consistency and active prompting methods (wang2022self, ; kojima2022large, ; diao2023active, ). In addition, yao2023tree proposes Tree-of-Thought, enabling models to explore multiple reasoning paths for solving complex problems, while besta2024graph introduces Graph-of-Thought, structuring LLM thoughts as graph vertices to enhance reasoning. We focus on CoT in this paper due to its widespread adoption and relative simplicity, while noting that our framework and methodology remain applicable to other reasoning approaches.
@@
Sensitivity and stability of explanations. In the vision domain,  dombrowski2019explanations ; heo2019fooling shows that the models are easy to be fooled with wrong explanations with correct predictions. In the language domain, to effectively evaluate the stability (situ2021learning, ) and faithfulness (madsen2024faith, ; chuang2024faith, ; lanham2023measuringfaith, ) of LLM explanations, as well as many works (tanneru2024quantifying, ; chen2023counter, ; ye2022unreliability, ) on quantifying uncertainty, counter-factual simulatability, and reliability of explanations. Compared with their work, MATCHA is the first evaluation framework for measuring the sensitivity of CoTs against token-level and embedding-level perturbations, where we provide a novel angle into the reasoning mechanism and robustness of different models.
@@
Adversarial attacks. Adversarial attacks against machine learning have been extensively explored for traditional application areas such as vision (pgd, ; fgsm, ; cw, ). Recent works have shown that LLMs are also susceptible to such attacks. In the LLM domain, gradient-based attacks like (gcg, ; zhu2023autodan, ; ji2024defending, ; liu2024automatic, ) show that carefully crafted text prefixes/suffixes allow a user to easily bypass LLM safety alignment training. priming shows that priming a model to start with an accepting phrase easily bypasses alignment. For embedding attacks,  schwinn2024soft reveals that subtle embedding-space manipulations can bypass safety alignment and revive unlearned behaviors, while xhonneux2024efficient shows that such continuous attacks significantly challenge alignment and require dedicated defenses. For other robustness definitions, zhou2024can explores techniques to enhance resilience against noisy rationales, xu2024preemptive designs a CoT attack using preemptive answers but does not assess true reasoning, and he2024using leverages CoT to improve robustness on challenging questions. MATCHA is orthogonal to these approaches as it focuses on examining the inconsistency between reasoning and answer with novel token and embedding-level perturbations.
@@
LLM evaluations. Existing studies evaluate LLMs on accuracy, robustness, and alignment with human values. Benchmarks like GLUE and SuperGLUE (glue, ; superglue, ) assess task performance, while HELM (helm, ) provides a holistic evaluation of safety, fairness, and efficiency. Quacer-B (quacer, ) introduces a certification framework for detecting LLM bias. On the other hand, LLM-as-a-judge is widely used (zhu2023judgelm, ; zheng2023judging, ; li2024generationjudge, ; huang2024empiricaljudge, ) in evaluating different capabilities of chat-boxes and LLMs. MATCHA can be used in parallel with these works as it focuses on the novel task of evaluating LLMs on the robustness of their reasoning using LLMs as judges.
@@ 3 Background
Adversarial attacks. Adversarial examples are a class of robustness attacks on neural networks (goodfellow2014explaining, ; kurakin2018adversarial, ). In image classification, we consider samples $\{(x_{i},y_{i})\}_{i=0}^{N}$ from an empirical distribution $\mathcal{D}$, where each image $x\in\mathbb{R}^{d}$ has a corresponding label $y\in\mathbb{R}^{k}$. A classifier $f$, parameterized by $\theta$, is trained to minimize a loss function $\mathcal{L}:\mathbb{R}^{k}\times\mathbb{R}^{k}\rightarrow\mathbb{R}$. Adversarial attacks seek a local point $x^{\prime}$ within a predefined adversarial region around $x$, i.e. $B_{p}(x,\epsilon_{p})=\{x^{\prime}\in\mathbb{R}^{d}:\|x^{\prime}-x\|_{p}\leq\epsilon_{p}\}$, such that the classification of $f$ changes ($f(x^{\prime})\neq f(x)=y$). $x^{\prime}$ and $x$ should be semantically identical, i.e. a human should determine them as being of the same class. A simple adversarial attack, introduced by goodfellow2014explaining , linearizes the loss function to compute perturbations that maximize the loss while adhering to adversarial constraints, requiring several backpropagation steps:
@@
We aim to generate the most probable pair of the answer $y_{p}$ and CoT $r_{p}$ from an LLM. We define $y_{p}$ as a preemptive answer, which means we generate the answer first then the reasoning, which poses a significant real-world threat (xu2024preemptive, ). When the model first produces an answer, we anchor the output, ensuring the reasoning step does not affect the answer. This allows us to explicitly control whether the reasoning aligns with the fixed correct answer (or intentionally diverges for perturbation purposes). In contrast, reasoning-first approaches intertwine reasoning with answering, making it harder to control one without affecting the other - we notice that if generating the reasoning first, then wrong reasoning will highly likely affect the answer to be wrong.
@@ 4 MATCHA: Misaligned Answer and Thought CHAin
In this section, we introduce a novel evaluation framework MATCHA that constructs perturbations that selectively degrade CoT quality while preserving the final answer. Unlike traditional perturbations that target model predictions directly, our goal is more nuanced: to expose discrepancies between correct answers and faulty reasoning, thereby revealing hidden failure modes in the reasoning process. Achieving this demands a carefully crafted optimization strategy. To be more specific, we present two complementary methods: token-level perturbations, which manipulate the input at the discrete level, and embedding-level perturbations, which operate directly in the continuous representation space. Each method presents its challenges and reveals different aspects of the model’s vulnerability.
4.1 Producing Correct-Answer Wrong-Reasoning Responses
@@
{mdframed}
[backgroundcolor=pink!20, linecolor=black] ICL ($n_{1}$): A list of (Q, CoT, A). {mdframed}[backgroundcolor=yellow!20, linecolor=black] Q ($n_{2}$): If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? {mdframed}[backgroundcolor=green!20, linecolor=black] A ($n_{3}$): The answer is 5. {mdframed}[backgroundcolor=blue!30, linecolor=black] CoT ($n_{4}$): Let’s think step by step. There are 3 cars at first. 2 more cars arrive. 3 + 2 = 5.
@@ 4.2 Token-Level Perturbations
{mdframed}
[backgroundcolor=yellow!20, linecolor=black] Q ($n_{2}$): If _ there are _3 cars_ in the _ parking lot and _ 2 more _ cars _ arrive, _ how many _ cars are _ in the _ parking lot?
@@ 4.3 Embedding-Level Perturbations
Embedding-level perturbations provide a powerful lens into LLM vulnerabilities beyond the discrete token space schwinn2024soft ; xhonneux2024efficient . Thus, we leverage embedding-level perturbations to subtly distort the model’s reasoning path while preserving answer correctness, exposing latent instabilities that are not easily revealed by token-level edits. The high-level idea is to perturb the embedding space of the input questions ($E_{c}[n_{1}:n_{2}]$) imperceptibly, such that it does not change the output token mapping, leading to the wrong reasoning yet correct answer. Similar to token-level perturbations, we generate the predicted answer $y_{p}$, $r_{p}$, and check whether $y_{p}$ equals $y_{gt}$. If so, we then generate target logits $q_{c},q_{a}$ using input embedding $E_{c}$. After that, we generate the perturbed logits $q^{\prime}_{c},q^{\prime}_{a}$ by passing in the perturbed embedding inputs for CoT and answer parts ($E^{\prime}_{c}$ and $E^{\prime}_{c}[:n_{3}]$). Given step-size $\alpha$, perturbation region $\epsilon$ (a percentage of the original embedding space), and using Eq. 3 to calculate $\mathcal{L}_{opt}$, we update the embedding by performing $l_{\infty}$ perturbation as follows:
@@
Then, we clamp $E_{tmp}$ into the interval of $[E_{c}-\epsilon,E_{c}+\epsilon]$ and only update only $E^{\prime}_{c}[n_{1}:n_{2}]$ part of the embedding with $E_{tmp}$, since we only perturb the embedding space of the question part. We continue the iterations for the perturbation with a budget of $j$ steps. After that, we generate the perturbed answer $y^{\prime}_{p}$ and CoT $r^{\prime}_{p}$ with perturbed input embedding $E^{\prime}_{c}$. To know whether the perturbation is successful, we follow the same procedure as the token-level perturbations. Algorithm 1 illustrates the embedding-level perturbation procedure of one unlabeled question.

### v2

# Robust Answers, Fragile Logic: Probing the Decoupling Hypothesis in LLM Reasoning

@@ Abstract
While Chain-of-Thought (CoT) prompting has become a cornerstone for complex reasoning in Large Language Models (LLMs), the faithfulness of the generated reasoning remains an open question. We investigate the Decoupling Hypothesis: that correct answers often mask fragile, post-hoc rationalizations that are not causally tied to the model’s prediction. To systematically verify this, we introduce MATCHA, a novel Answer-Conditioned Probing framework. Unlike standard evaluations that focus on final output accuracy, MATCHA isolates the reasoning phase by conditioning generation on the model’s predicted answer, allowing us to stress-test the stability of the rationale itself. Our experiments reveal a critical vulnerability: under imperceptible input perturbations, LLMs frequently maintain the correct answer while generating inconsistent or nonsensical reasoning - effectively being “Right for the Wrong Reasons”. Using LLM judges to quantify this robustness gap, we find that multi-step and commonsense tasks are significantly more susceptible to this decoupling than logical tasks. Furthermore, we demonstrate that adversarial examples generated by MATCHA transfer non-trivially to black-box models. Our findings expose the illusion of CoT robustness and underscore the need for future architectures that enforce genuine answer-reasoning consistency rather than mere surface-level accuracy.
@@ 1 Introduction
Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023), Llama-3 (Dubey et al., 2024b), and DeepSeek-R1 (Guo et al., 2025a) have demonstrated remarkable capabilities in complex reasoning through Chain-of-Thought (CoT) prompting (Wei et al., 2022a; b). The prevailing assumption is that these intermediate steps reflect a faithful causal process: the model “thinks”, and therefore it “answers”. However, we challenge the validity of this causal link. We propose the Decoupling Hypothesis: the generation of reasoning paths and the selection of the answer are not strictly causally bound, but rather loosely correlated. While CoT is widely deployed for explanation in education Grassucci et al. (2025) and healthcare Ali (2024), current evaluations fail to distinguish between a model that reasons to reach an answer and one that merely rationalizes a predetermined decision.
To rigorously disentangle these two processes, we introduce a novel Answer-Conditioned Probing framework. Unlike standard adversarial attacks that target the final output label, our unique setting isolates the reasoning phase by conditioning the rationale generation on the model’s own predicted answer. This allows us to simulate a critical cognitive failure mode: akin to a student who guesses the correct answer by intuition but hallucinates a derivation to justify it. Within this framework, we explicitly attack the rationale generation while the answer remains committed. We demonstrate that imperceptible perturbations can collapse the logical consistency of the reasoning chain without flipping the answer. This phenomenon reveals that models can be “Right for the Wrong Reasons”, creating a false sense of security where robust answers mask fragile, unfaithful reasoning.
@@
In this paper, we show the fragility of LLM CoT, as exemplified in Figure 1, where a perturbed question leads to the wrong reasoning yet correct answer (DeepSeek-R1-7 on GSM8K). We adopt an answer-first, then reasoning framework to isolate failures in reasoning while keeping the answer fixed. This setup not only exposes reasoning-specific vulnerabilities but also reflects real-world scenarios where LLM reasoning is often generated post hoc to justify already-decided outputs Jeyasothy et al. (2023); Xu et al. (2024). To this end, we introduce MATCHA, a novel algorithm for finding small input changes that cause inconsistent behavior in CoT reasoning. We use MATCHA to create metrics for evaluating the robustness of an LLM’s CoT for different models. MATCHA can be instantiated as either a token-level or embedding-level perturbation. Our primary focus is, token-level perturbations as they are more realistic; however, they are challenging to create due to the need to introduce minimal, semantically plausible edits that selectively disrupt reasoning without altering answer correctness. To achieve this, we first randomly insert tokens into the original question. Then, we identify the inserted tokens that have the largest influence on the reasoning process while ensuring that answer-related tokens remain unchanged. By leveraging gradient-based importance ranking, we selectively replace tokens that maximize reasoning perturbation while minimizing changes to the original question. MATCHA as an embedding-level perturbation perturbs the embedding space of input questions imperceptibly, ensuring that the token sequence remains unchanged while disrupting the reasoning process. By optimizing a loss that maximizes reasoning divergence while preserving the correct answer, the perturbation subtly shifts the internal representations, leading to incorrect reasoning without altering the final answer.
@@
  • To our knowledge, we are the first to systematically probe this Decoupling Hypothesis. By leveraging our Answer-Conditioned framework, we quantify the critical misalignment between reasoning stability and answer accuracy. We introduce a novel evaluation framework MATCHA (Misaligned Answer and Thought CHAin) which works at both the token- and embedding-level.
@@ 2 Related Work
Prompt-based Reasoning. Chain-of-Thought was first introduced by Wei et al. (2022a), showing that inducing intermediate reasoning steps significantly improves the reasoning ability of LLMs, which is widely regarded as an important form of NLEs. Many works have further expanded on Chain-of-Thought (CoT) using self-consistency and active prompting methods (Wang et al., 2022; Kojima et al., 2022a; Diao et al., 2023). In addition, Yao et al. (2023) proposes Tree-of-Thought, enabling models to explore multiple reasoning paths for solving complex problems, while Besta et al. (2024) introduces Graph-of-Thought, structuring LLM thoughts as graph vertices to enhance reasoning. We focus on CoT in this paper due to its widespread adoption and relative simplicity, while noting that our framework and methodology remain applicable to other reasoning approaches.
@@
Sensitivity and stability of explanations. In the vision domain,  Dombrowski et al. (2019); Heo et al. (2019) shows that the models are easy to be fooled with wrong explanations with correct predictions. In the language domain, to effectively evaluate the stability (Situ et al., 2021) and faithfulness (Madsen et al., 2024; Chuang et al., 2024; Lanham et al., 2023) of LLM explanations, as well as many works (Tanneru et al., 2024; Chen et al., 2023; Ye and Durrett, 2022) on quantifying uncertainty, counter-factual simulatability, and reliability of explanations. Compared with their work, MATCHA is the first evaluation framework for measuring the sensitivity of CoTs against token-level and embedding-level perturbations, where we provide a novel angle into the reasoning mechanism and robustness of different models.
@@
Adversarial attacks. Adversarial attacks against machine learning have been extensively explored for traditional application areas such as vision (Madry et al., 2018; Goodfellow et al., 2014a; Carlini and Wagner, 2017). Recent works have shown that LLMs are also susceptible to such attacks. In the LLM domain, gradient-based attacks like (Zou et al., 2023; Zhu et al., 2023b; Ji et al., 2024; Liu et al., 2024) show that carefully crafted text prefixes/suffixes allow a user to easily bypass LLM safety alignment training. Vega et al. (2023) shows that priming a model to start with an accepting phrase easily bypasses alignment. For embedding attacks,  Schwinn et al. (2024) reveals that subtle embedding-space manipulations can bypass safety alignment and revive unlearned behaviors, while Xhonneux et al. (2024) shows that such continuous attacks significantly challenge alignment and require dedicated defenses. For other robustness definitions, Zhou et al. (2024) explores techniques to enhance resilience against noisy rationales, Xu et al. (2024) designs a CoT attack using preemptive answers but does not assess true reasoning, and He et al. (2024) leverages CoT to improve robustness on challenging questions. MATCHA is orthogonal to these approaches as it focuses on examining the inconsistency between reasoning and answer with novel token and embedding-level perturbations.
@@
LLM evaluations. Existing studies evaluate LLMs on accuracy, robustness, and alignment with human values. Benchmarks like GLUE and SuperGLUE (Wang, 2018; Sarlin et al., 2020) assess task performance, while HELM (Liang et al., 2022) provides a holistic evaluation of safety, fairness, and efficiency. Quacer-B (Chaudhary et al., 2024) introduces a certification framework for detecting LLM bias. On the other hand, LLM-as-a-judge is widely used (Zhu et al., 2023a; Zheng et al., 2023; Li et al., 2024; Huang et al., 2024) in evaluating different capabilities of chat-boxes and LLMs. MATCHA can be used in parallel with these works as it focuses on the novel task of evaluating LLMs on the robustness of their reasoning using LLMs as judges.
@@ 3 Background
Adversarial attacks. Adversarial examples are a class of robustness attacks on neural networks (Goodfellow et al., 2014b; Kurakin et al., 2018). In image classification, we consider samples $\{(x_{i},y_{i})\}_{i=0}^{N}$ from an empirical distribution $\mathcal{D}$, where each image $x\in\mathbb{R}^{d}$ has a corresponding label $y\in\mathbb{R}^{k}$. A classifier $f$, parameterized by $\theta$, is trained to minimize a loss function $\mathcal{L}:\mathbb{R}^{k}\times\mathbb{R}^{k}\rightarrow\mathbb{R}$. Adversarial attacks seek a local point $x^{\prime}$ within a predefined adversarial region around $x$, i.e. $B_{p}(x,\epsilon_{p})=\{x^{\prime}\in\mathbb{R}^{d}:\|x^{\prime}-x\|_{p}\leq\epsilon_{p}\}$, such that the classification of $f$ changes ($f(x^{\prime})\neq f(x)=y$). $x^{\prime}$ and $x$ should be semantically identical, i.e. a human should determine them as being of the same class. A simple adversarial attack, introduced by Goodfellow et al. (2014b), linearizes the loss function to compute perturbations that maximize the loss while adhering to adversarial constraints, requiring several backpropagation steps:
@@
We aim to generate the most probable pair of an answer $y_{p}$ and a Chain-of-Thought $r_{p}$ from an LLM. We define $y_{p}$ as a preemptive answer, meaning that the answer is generated before the reasoning. This design choice allows us to rigorously isolate the robustness of the reasoning process via an Answer-Conditioned strategy. In standard reasoning-first approaches, the answer is causally downstream of the rationale; as a result, perturbations that degrade the reasoning often propagate to the final prediction. This coupling introduces a confounding factor that obscures the specific phenomenon of “Right for the Wrong Reasons”. By anchoring the model’s predicted answer first, we explicitly decouple decision-making from rationalization. This enables us to control the decision boundary and stress-test the reasoning in isolation, ensuring that any observed fragility reflects properties of the reasoning structure itself rather than being a consequence of answer changes.
@@ 4 MATCHA: Misaligned Answer and Thought CHAin
In this section, we introduce a novel evaluation framework MATCHA that constructs perturbations that selectively degrade CoT quality while preserving the generated correct answer. Unlike traditional perturbations that target model predictions directly, our goal is more nuanced: to expose discrepancies between correct answers and faulty reasoning, thereby revealing hidden failure modes in the reasoning process. Achieving this demands a carefully crafted optimization strategy. To be more specific, we present two complementary methods: token-level perturbations, which manipulate the input at the discrete level, and embedding-level perturbations, which operate directly in the continuous representation space. Each method presents its challenges and reveals different aspects of the model’s vulnerability.
4.1 Answer-Conditioned Probing
@@
ICL ($n_{1}$): A list of (Q, CoT, A). Q ($n_{2}$): If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? A ($n_{3}$): The answer is 5. CoT ($n_{4}$): Let’s think step by step. There are 3 cars at first. 2 more cars arrive. 3 + 2 = 5.
@@ 4.2 Token-Level Perturbations
Q ($n_{2}$): If _ there are _3 cars_ in the _ parking lot and _ 2 more _ cars _ arrive, _ how many _ cars are _ in the _ parking lot?
@@ 4.3 Embedding-Level Perturbations
Embedding-level perturbations provide a powerful lens into LLM vulnerabilities beyond the discrete token space Schwinn et al. (2024); Xhonneux et al. (2024). Thus, we leverage embedding-level perturbations to subtly distort the model’s reasoning path while preserving answer correctness, exposing latent instabilities that are not easily revealed by token-level edits. The high-level idea is to perturb the embedding space of the input questions ($E_{c}[n_{1}:n_{2}]$) imperceptibly, such that it does not change the output token mapping, leading to the wrong reasoning yet correct answer. Similar to token-level perturbations, we generate the predicted answer $y_{p}$, $r_{p}$, and check whether $y_{p}$ equals $y_{gt}$. If so, we then generate target logits $q_{c},q_{a}$ using input embedding $E_{c}$. After that, we generate the perturbed logits $q^{\prime}_{c},q^{\prime}_{a}$ by passing in the perturbed embedding inputs for CoT and answer parts ($E^{\prime}_{c}$ and $E^{\prime}_{c}[:n_{3}]$). Given step-size $\alpha$, perturbation region $\epsilon$ (a percentage of the original embedding space), and using Eq. 3 to calculate $\mathcal{L}_{opt}$, we update the embedding by performing $l_{\infty}$ perturbation as follows:
@@
Then, we clamp $E_{tmp}$ into the interval of $E_{c}-\epsilon,E_{c}+\epsilon]$ and only update only $E^{\prime}_{c}[n_{1}:n_{2}]$ part of the embedding with $E_{tmp}$, since we only perturb the embedding space of the question part. We continue the iterations for the perturbation with a budget of $j$ steps. After that, we generate the perturbed answer $y^{\prime}_{p}$ and CoT $r^{\prime}_{p}$ with perturbed input embedding $E^{\prime}_{c}$. To know whether the perturbation is successful, we follow the same procedure as the token-level perturbations. Algorithm 1 illustrates the embedding-level perturbation procedure of one unlabeled question.